OpenStudy (anonymous):

Elena buys several shares of stock at $28 per share and sells them at $24.50 per share. What is the percent of decrease in the price of the stock?

OpenStudy (anonymous):

it would be a 12.5 % decrease

OpenStudy (lilg132):

12.5 percent

OpenStudy (anonymous):

thank you guys sooo much i am horrible at algebra..lol

OpenStudy (lilg132):

28 / 100 = 0.28 (that is 1 percent of 28) 28 - 24.5 = 3.5 is the difference 3.5 / 0.28 = 12.5 percent

OpenStudy (lilg132):

dont forget to click good answer :p

OpenStudy (anonymous):

:)